A body travel the first 1/2 of the total distance with velocity v1 & second 1/2 with v2. Calculate the A.V.?

Grade:11

1 Answers

Vikas TU
7 years ago
Average velocity = total displacement/total time
for first half,
t1 = x/v1
for second half,
t2 = x/v2
total time = t1 + t2
avg. velocity = 2x/(t1+t2) = > 2x/(x/v1 + x/v2) = > 2/(1/v1 + 1/v2) m/s